### v1

# Optimal Dynamic Fees in Automated Market Makers

Abstract
Automated Market Makers (AMMs) are emerging as a popular decentralised trading platform. In this work, we determine the optimal dynamic fees in a constant function market maker. We find approximate closed-form solutions to the control problem and study the optimal fee structure. We find that there are two distinct fee regimes: one in which the AMM imposes higher fees to deter arbitrageurs, and another where fees are lowered to increase volatility and attract noise traders. Our results also show that dynamic fees that are linear in inventory and are sensitive to changes in the external price are a good approximation of the optimal fee structure and thus constitute suitable candidates when designing fees for AMMs.
Keywords: decentralized finance, automated market makers, optimal fees, arbitrageurs, noise trading.
Mathematics Subject Classification (2020): 93E20, 91B70.
JEL Classification: C61, G23, D53.
1 Introduction
In recent years, automated market makers (AMMs) have gained popularity as a key innovation in financial technology, fundamentally changing the way liquidity is provided and exchanged within decentralized finance (DeFi) ecosystems. An AMM is a decentralized venue, where liquidity provision and liquidity taking follow rules that are encoded in smart contracts. As of 2025, the most popular AMM is Uniswap, which began gaining widespread adoption with its v2 version and has now reached v4; see Adams et al. (2020, 2021, 2024) for the white papers on versions v2 to v4, and Bachu et al. (2025) for a detailed overview of the v4 protocol.
The early works of Capponi and Jia (2021), Bartoletti et al. (2021), Angeris and Chitra (2020), Angeris et al. (2021), and Angeris et al. (2023) first explain how these exchanges operate and introduce the problem of liquidity provision in AMMs. As shown by Cartea et al. (2023, 2024b) and Milionis et al. (2022), liquidity provision in the absence of fees leads to losses for liquidity providers. This issue, known as impermanent loss (IL), occurs because liquidity provision in an AMM is passive; when the prices of risky assets move, the liquidity providers trade at worse than market prices.
To address the impermanent loss problem, a significant part of the research has focused on improving the design of AMMs to make liquidity provision less exposed to arbitrageurs. Some works in this direction are Adams et al. (2025), where they set up an auction-based AMM in which LPs are rewarded by a so-called manager of the pool who can profit from arbitrage opportunities; Goyal et al. (2023), where they translate LPs’ beliefs about future asset valuations into an optimal choice of a trading function; Cartea et al. (2024a), where they introduce a more general version of constant function market makers (CFMs) called decentralized liquidity pools (DLPs), in which liquidity providers prevent predictable losses by choosing particular quote functions, impact functions, and monitoring external information; and Bergault et al. (2024b), where they design an oracle-based AMM in which price discovery does not rely solely on liquidity takers, and the pricing function incorporates external information about the current market price of the risky asset. Recently Capponi and Zhu (2025) addressed the problem of IL by determining an optimal exit strategy for the liquidity provider. Another interesting contribution is the work of Fukasawa et al. (2023), where the authors hedge impermanent loss in a constant function market maker using a weighted variance swap.
As shown by Milionis et al. (2024), impermanent loss can be partially or fully hedged through the collection of fees. The problem of determining optimal fees for AMMs has been an important issue since the early stages of research in this area. Evans et al. (2021) focused on how to find optimal fees for geometric mean market makers (G3Ms) while Fritsch (2021) studied how to determine optimal fees in a setting of multiple liquidity pools competing with each other for trade volume. However, both works only address the problem of finding constant static fees. Sabate-Vidales and Šiška (2023) address the impermanent loss problem by setting dynamic fees although no optimality is established. Hasbrouck et al. (2022) found how higher fees on decentralised exchanges (DEX) can increase trading volume because higher fees attract more liquidity providers and boost the inventory in the DEX. Recently, Aqsha et al. (2025) solve a stochastic leader-follower game between the AMM and the liquidity providers in the venue. They find that under the equilibrium contract, the liquidity providers have incentives to add liquidity to the pool only when higher liquidity on average attracts more noise trading. In their work, the fee that the venue charges is proportional to the trade size.
In this study, we address the problem of determining optimal dynamic fees for AMMs. As argued in Cao et al. (2023), the fees should adjust dynamically in order to reflect market conditions. With the recent introduction of Uniswap v4, the implementation of dynamic fees is now possible. Setting optimal fees is crucial because fees are the primary incentive for liquidity providers to engage with the AMM as they help offset losses from impermanent loss. As a result, a platform that optimizes fee collection can better compensate liquidity providers and, in turn, strengthen its overall market position. Inspired by the work on market making by Avellaneda and Stoikov (2008), [^1] For a comprehensive exposition of the market making problem and its generalisation see the monographs by Cartea et al. (2015) and Gueant (2016). we formulate a stochastic control problem in which the venue aims to maximize the total fees collected in a liquidity pool consisting of two assets: a risky asset $Y$ and a riskless asset $X$. The control processes are $\mathfrak{p}$ and $\mathfrak{m}$, representing the fees charged for selling and buying asset $Y$, respectively. [^2] We choose $\mathfrak{p}$ for plus and $\mathfrak{m}$ for minus because the quantity of asset $Y$ in the pool increases after a trade that pays $\mathfrak{p}$ and decreases after a trade that pays $\mathfrak{m}$.
These fees influence the exchange rate faced by liquidity takers: buying $Y$ requires paying more of asset $X$, while selling $Y$ yields less of asset $X$. This, in turn, affects the order flow. We model sell and buy orders using two point processes,$\{N^{+,\mathfrak{p}}_{t}\}_{t\in[0,T]}$ and $\{N^{-,\mathfrak{m}}_{t}\}_{t\in[0,T]}$, with controlled stochastic intensities of the form:
$\lambda_{t}^{+,\mathfrak{p}}$ $:=\lambda^{+}\exp\left(k\left(Z_{+}^{\mathfrak{p}_{t}}(y_{t-}^{\mathfrak{p},\mathfrak{m}})-(S_{t}+\zeta)\right)\Delta^{+}(y_{t-}^{\mathfrak{p},\mathfrak{m}})\right)\mathbbm{1}_{\{y_{t-}^{\mathfrak{p},\mathfrak{m}}<\overline{y}\}},$
$\lambda_{t}^{-,\mathfrak{m}}$ $:=\lambda^{-}\exp\left(-k\left(Z_{-}^{\mathfrak{m}_{t}}(y_{t-}^{\mathfrak{p},\mathfrak{m}})-(S_{t}-\zeta)\right)\Delta^{-}(y_{t-}^{\mathfrak{p},\mathfrak{m}})\right)\mathbbm{1}_{\{y_{t-}^{\mathfrak{p},\mathfrak{m}}>\underline{y}\}}.$
Here, $S_{t}$ denotes the price of asset $Y$ in terms of asset $X$ on a centralized venue (e.g., Kraken, Binance), while $\Delta^{-}(y_{t-}^{\mathfrak{p},\mathfrak{m}})$ and $\Delta^{+}(y_{t-}^{\mathfrak{p},\mathfrak{m}})$ represent the amounts that a liquidity taker trades when buying and selling asset $Y$, respectively. The control problem we wish to solve is, in its full generality, intractable. Therefore, we study two simplified cases using different approximations. First, we assume that the external price $S_{t}$ remains constant at $S_{0}$. Under this assumption, we derive in Theorem 4.1 the optimal fee policies and the corresponding value function. Second, we approximate the exponential terms in the Hamilton–Jacobi–Bellman (HJB) equation using a linear-quadratic expansion. This leads to an alternative tractable formulation from which we derive the optimal fees and value function, as stated in Theorem 5.2. In both approximations, we obtain the same insight: the venue finds it optimal to adopts two distinct fee regimes; one in which the AMM imposes higher fees to deter arbitrageurs, and another where fees are lowered to increase volatility and attract noise traders. Figures 1(a) and 1(b) illustrate the optimal buying and selling fees under the first (left) and second (right) approximations, respectively, with the two fee patterns highlighted.
(a) Optimal fees for selling $\mathfrak{p}_{t}^{*}$ (solid line) and for buying $\mathfrak{m}_{t}^{*}$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool with the approximation that takes $S_{t}=S_{0}$ for $t\in[0,T]$.
(b) Optimal fees for selling $\mathfrak{p}_{t}^{*}$ (solid line) and for buying $\mathfrak{m}_{t}^{*}$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool with the linear quadratic approximation of the stochastic intensities.
Finally, we run simulations to compare the performance of the optimal fee policies against alternative strategies. In the case of a constant external price, we find that the optimal fees outperform constant fee strategies by a significant margin. Furthermore, a linear approximation of the optimal policies proves to be sufficiently accurate, yielding nearly identical revenue outcomes. Our results suggest that the optimal fees balance two competing objectives: setting fees high enough to maximize per-trade revenue while simultaneously increasing the quadratic variation of the marginal price, thereby attracting increased noise trading. A similar pattern emerges under the second approximation, where the revenue generated by the optimal fees also exceeds that of the alternative strategies.
The remainder of the paper is organised as follows. Section 2 introduces the model that we use, sets up the control problem and the performance criterion that the venue wants to maximize. In Section 3 we compute the optimal policies. Sections 4 and 5 are dedicated to finding the value functions for the first and second approximation, respectively. Finally, in Sections 4.1 and 5.1 we carry out simulations to illustrate our theoretical results.
2 The model
We study an automated market maker (AMM) with a finite time horizon $T>0$. More precisely, we consider a constant function market (CFM) pool inside the reference AMM with a riskless asset $X$ and a risky asset $Y$. Let $00pt>$ 0 denote the depth of the pool and let $f:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ be the trading function, which is strictly increasing and twice differentiable in both arguments. The trading condition is $f(x,y)=0pt$, where $x$ and $y$ denote the amounts in assets $X$ and $Y$, respectively. We can rewrite this condition using the level function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ satisfying $f(\varphi(y),y)=0pt$. The latter is then automatically differentiable and decreasing, and we assume in addition that it is convex in order to exclude roundtrip arbitrage, see Cartea et al. (2024b). These assumptions are satisfied by the most popular CFMs such as constant product markets (CPMs) where $f(x,y)=xy$ and $\varphi(y)=0pt/y$.
We assume that quantity of asset $Y$ in the pool takes values in a finite grid given by
$\{y^{-N}:=\underline{y},\dots,y^{0},\dots,y^{N}:=\overline{y}\},$
where $\underline{y}$ and $\overline{y}$ can be interpreted as reserve constraints satisfying $0<\underline{y}<y^{0}<\overline{y}<0pt$.
Consequently, the quantity of asset $X$ in the pool takes values in the grid
$\{x^{-N}:=\varphi(y^{-N}),\dots,x^{0}:=\varphi(y^{0}),\dots,x^{N}:=\varphi(y^{N})\}.$
Note that the monotonicity of the grids for the two assets are opposite to each other: the grid for asset $Y$ is increasing, whereas the grid for asset $X$ is decreasing.
There are three basic exchange rates for the price of asset $Y$ in units of asset $X$ in terms of the quantity of asset $Y$ in the pool that we use throughout the paper. To fix notation, let $y^{i}$ denote the quantity of asset $Y$ in the pool just before the trade of a liquidity taker.
  1. (a)
    The marginal exchange rate describes the price of an infinitesimal trade and is given by
    $Z(y^{i}):=-\varphi^{\prime}(y^{i}).$
  2. (b)
    The exchange rate for buying (taking out of the pool) $\Delta^{-}(y^{i}):=y^{i}-y^{i-1}$ units of asset $Y$ is given by
    $Z_{-}(y^{i}):=\frac{\varphi(y^{i-1})-\varphi(y^{i})}{\Delta^{-}(y^{i})}.$
    It takes values in the grid
    $\{Z_{-}^{-N+1}:=Z_{-}(y^{-N+1}),\dots,Z_{-}^{0}:=Z_{-}(y^{0}),\dots,Z_{-}^{N}:=Z_{-}(y^{N})\}.$
  3. (c)
    The exchange rate for selling (depositing in the pool) $\Delta^{+}(y^{i}):=y^{i+1}-y^{i}$ units of asset $Y$ is given by
    $Z_{+}(y^{i}):=\frac{\varphi(y^{i})-\varphi(y^{i+1})}{\Delta^{+}(y^{i})}.$
    It takes values in the grid
    $\{Z_{+}^{-N}:=Z_{+}(y^{-N}),\dots,Z_{+}^{0}:=Z_{+}(y^{0}),\dots,Z_{+}^{N-1}:=Z_{+}(y^{N-1})\}.$
Note that exchange rates for buying and selling satisfy the identity
$Z_{-}^{i}=Z_{+}^{i-1},\quad i\in\{-N+1,\dots,0,\dots N\}.$
We now add proportional transaction cost to the trading mechanism, and model the fees for buying and selling by the functions $\mathfrak{m}:\{y^{-N},\ldots,y^{N}\}\to\mathbb{R}$ and $\mathfrak{p}:\{y^{-N},\ldots,y^{N}\}\to\mathbb{R}$, respectively. [^3] Although one would expect that $\mathfrak{m},\mathfrak{p}\in[0,1]$, we do not impose this condition since it turns out to be optimal to sometimes charge negative fees, meaning that the venues pays liquidity takers to trade in a given direction. The fees are collected in units of the riskless asset $X$ and are deposited outside of the pool. The corresponding (fee-depending) exchange rates can then be described in the following way, where $y^{i}$ denotes the quantity of asset $Y$ in the pool just before the trade.
  1. (a)
    The exchange rate with fee structure $\mathfrak{m}$ for buying (taking out of the pool) $\Delta^{-}(y^{i}):=y^{i}-y^{i-1}$ units of asset $Y$ is given by
    $Z_{-}^{\mathfrak{m}}(y^{i}):=(1+\mathfrak{m}(y^{i}))Z_{-}(y^{i}).$
  2. (b)
    The exchange rate with fee structure $\mathfrak{p}$ for selling (depositing in the pool) $\Delta^{+}(y^{i}):=y^{i+1}-y^{i}$ units of asset $y$ is given by
    $Z_{+}^{\mathfrak{p}}(y^{i}):=(1-\mathfrak{p}(y^{i}))Z_{+}(y^{i}).$
In this paper the AMM mechanism works independently of the fee structure because the fees charged by the venue go to the venue’s cash account to be redistributed at a later stage. In the pool, trading is carried out via the rates $Z_{+}$ and $Z_{-}$ so that the depth $0pt$ does not change. However, from the point of view of the liquidity taker, the exchange rates include fees.
Our model aims to study how the AMM sets fees dynamically in order to maximize revenue coming from the order flow. We fix a probability space $(\Omega,\mathcal{F},\mathbb{F}=\{\mathcal{F}_{t}\}_{t\in[0,T]},\mathbb{P}^{\mathfrak{p},\mathfrak{m}})$ supporting all the processes we use below. [^4] A rigorous definition of the probability space would involve a weak formulation of the control problem because the controls affect the intensities of the jump processes that inform the filtration. We leave this rigorous treatment out of the paper for readability. See Section 3.1 in Barucci et al. (2025) for a formal treatment of this issue. We model one representative liquidity taker (LT) and assume that their buy and sell order arrivals are described by controlled point processes $\{N^{-,\mathfrak{m}}_{t}\}_{t\in[0,T]}$ and $\{N^{+,\mathfrak{p}}_{t}\}_{t\in[0,T]}$, respectively, where the fee structure processes $\{\mathfrak{p}_{t}\}_{t\in[0,T]}$ and $\{\mathfrak{m}_{t}\}_{t\in[0,T]}$ are assumed to be predictable. In this work, the depth of the pool $0pt$ is fixed throughout $[0,T]$, that is, $T$ is small enough so that LPs do not add or remove liquidity from the pool. This allows us to study liquidity taking and optimal fees in isolation.
The controls will operate in such a way that the difference $N^{+,\mathfrak{p}}-N^{-,\mathfrak{m}}$ takes values in $\{-N,\ldots,N\}$. The quantity of asset $Y$ at time $t\in[0,T]$ is then given by
$Y_{t}^{\mathfrak{p},\mathfrak{m}}:=y^{N^{+,\mathfrak{p}}_{t}-N^{-,\mathfrak{m}}_{t}}.$
The controlled intensities of $\{N^{-,\mathfrak{m}}_{t}\}_{t\in[0,T]}$ and $\{N^{+,\mathfrak{p}}_{t}\}_{t\in[0,T]}$ are given by
$\lambda_{t}^{-,\mathfrak{m}}$ $:=\lambda^{-}\exp{\left(-k(Z_{-}^{\mathfrak{m}_{t}}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})-(S_{t}-\zeta))\Delta^{-}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})\right)}\mathbbm{1}_{\{Y_{t-}^{\mathfrak{p},\mathfrak{m}}>\underline{y}\}},$
$\lambda_{t}^{+,\mathfrak{p}}$ $:=\lambda^{+}\exp{\left(k(Z_{+}^{\mathfrak{p}_{t}}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})-(S_{t}+\zeta))\Delta^{+}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})\right)}\mathbbm{1}_{\{Y_{t-}^{\mathfrak{p},\mathfrak{m}}<\overline{y}\}}.$
Here, $\lambda^{-}$ and $\lambda^{+}$ are some baseline intensities, $k>0$ is an exponential decay parameter (similar to that in Avellaneda and Stoikov (2008)), $\{S_{t}\}_{t\in[0,T]}$ denotes the midprice in a centralised exchange (outside of the pool) of asset $Y$ in terms of the riskless asset $X$, the so-called oracle price, and $\zeta>0$ is the corresponding half-spread in the centralised exchange. We assume that $\{S_{t}\}_{t\in[0,T]}$ has dynamics given by
$S_{t}=S_{0}+\sigma W_{t},$
where $\{W_{t}\}_{\{t\in[0,T]\}}$ is a Brownian motion.
The controlled stochastic intensities $\lambda_{t}^{-,\mathfrak{m}}$ and $\lambda_{t}^{+,\mathfrak{p}}$ aim to capture the following stylised fact: when $Z_{-}^{\mathfrak{m}_{t}}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})<S_{t}-\zeta$, buying $\Delta^{-}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})$ units to the pool and selling the same amount from the external venue is profitable to arbitrageurs. Thus, all else being equal, the buy intensity $\lambda_{t}^{+,\mathfrak{p}}$ increases. Similarly, when $Z_{+}^{\mathfrak{p}_{t}}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})>S_{t}+\zeta$, selling $\Delta^{+}(Y_{t-}^{\mathfrak{p},\mathfrak{m}})$ units to the pool and buying the same amount from the external venue is profitable to arbitrageurs, and the sell intensity $\lambda_{t}^{+,\mathfrak{p}}$ increases. In order to make the notation simpler, we incorporate $e^{-k\zeta}$ into $\lambda^{+}$ and $\lambda^{-}$ and formally set $\zeta:=0$.
The cumulative fees $\{\mathfrak{C}_{t}^{\mathfrak{p},\mathfrak{m}}\}_{t\in[0,T]}$ collected by the AMM are in turn given by
$\mathfrak{C}_{t}^{\mathfrak{p},\mathfrak{m}}=\int_{0}^{t}\mathfrak{p}_{u}Z_{+}^{\mathfrak{p}_{u}}(Y_{u^{-}}^{\mathfrak{p},\mathfrak{m}})\Delta^{+}(Y_{u^{-}}^{\mathfrak{p},\mathfrak{m}})\,\mathrm{d}N^{+,\mathfrak{p}}_{u}+\int_{0}^{t}\mathfrak{m}_{u}Z_{-}^{\mathfrak{m}_{u}}(Y_{u^{-}}^{\mathfrak{p},\mathfrak{m}})\Delta^{-}(Y_{u^{-}}^{\mathfrak{p},\mathfrak{m}}))\,\mathrm{d}N^{-,\mathfrak{m}}_{u},\quad\quad t\in[0,T].$
The performance criterion of the venue is given by
$J(\mathfrak{m},\mathfrak{p}):=\mathbb{E}\left[\mathfrak{C}_{T}^{\mathfrak{p},\mathfrak{m}}-\int_{0}^{T}P(Y_{t}^{\mathfrak{p},\mathfrak{m}},S_{t})\,\mathrm{d}t\right],$
where $P$ is a penalty function. An example of the function $P$ that we consider below is the quadratic distance between the oracle price $S_{t}$ and the instantaneous exchange rate $Z(y_{t})$, i.e.,
$P(Y_{t}^{\mathfrak{p},\mathfrak{m}},S_{t})=\phi(Z(Y_{t}^{\mathfrak{p},\mathfrak{m}})-S_{t})^{2}.$
Note, however, that $P$ is not needed to add concavity to the problem, because the term $\mathfrak{C}_{T}$ itself is concave in the controls. Our main results and insights are obtained when $P=0$.
3 Characterization of the value function
The AMM seeks to solve the control problem
$v(t,\mathfrak{c},y,s):=\sup_{(\mathfrak{m},\mathfrak{p})\in\mathcal{A}_{t}}v^{(\mathfrak{m},\mathfrak{p})}(t,\mathfrak{c},y,s),$
where $\mathcal{A}_{t}$ denotes the set of all $\mathbb{F}$-predictable and bounded fee structure processes $(\mathfrak{m}_{u},\mathfrak{p}_{u})_{\{t\leq u\leq T\}}$ and the conditional performance criterion is given by
$v^{(\mathfrak{m},\mathfrak{p})}(t,\mathfrak{c},y,s):=\mathbb{E}_{(t,\mathfrak{c},y,s)}\left[\mathfrak{C}_{T}^{(t,\mathfrak{c},y,s,\mathfrak{m},\mathfrak{p})}-\int_{t}^{T}P(Y^{(t,\mathfrak{c},y,s,\mathfrak{m},\mathfrak{p})}_{u},S_{u}^{(t,\mathfrak{c},y,s)})\,\mathrm{d}u\right].$
Here, $\{\mathfrak{C}_{u}^{(t,\mathfrak{c},y,s,\mathfrak{m},\mathfrak{p})}\}_{u\in[t,T]}$, $\{Y_{u}^{(t,\mathfrak{c},y,s,\mathfrak{m},\mathfrak{p})}\}_{u\in[t,T]}$ and $\{S_{u}^{(t,\mathfrak{c},y,s)}\}_{u\in[t,T]}$ denote the (controlled) processes $\mathfrak{C}$, $y$ and $S$ restarted at time $t$ with initial value $\mathfrak{c}$, $y$ and $s$, respectively. From the dynamic programming principle, we determine that the Hamilton-Jacobi-Bellman (HJB) equation satisfied by the value function is:
$\frac{\partial}{\partial t}v(t,y,\mathfrak{c},s)+\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial s^{2}}v(t,y,\mathfrak{c},s)-P(y,s)$
$+\left(\lambda^{+}e^{k(Z_{+}(y)-s)\Delta^{+}(y)}\sup_{\mathfrak{p}\in\mathbb{R}}e^{-k\mathfrak{p}Z_{+}(y)\Delta^{+}(y)}\left[v(t,y+\Delta^{+}(y),\mathfrak{c}+\mathfrak{p}Z_{+}(y)\Delta^{+}(y),s)-v(t,\mathfrak{c},y,s)\right]\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.1)
$+\left(\lambda^{-}e^{-k(Z_{-}(y)-s)\Delta^{-}(y)}\sup_{\mathfrak{m}\in\mathbb{R}}e^{-k\mathfrak{m}Z_{-}(y)\Delta^{-}(y)}\left[v(t,y-\Delta^{-}(y),\mathfrak{c}+\mathfrak{m}Z_{-}(y)\Delta^{-}(y),s)-v(t,\mathfrak{c},y,s)\right]\right)\mathbbm{1}_{\{y>\underline{y}\}}=0,$
with terminal condition $v(T,y,\mathfrak{c},s)=\mathfrak{c}$. The terms in this HJB equation have an intuitive interpretation. The terms
$v(t,y+\Delta^{+}(y),\mathfrak{c}+\mathfrak{p}Z_{+}(y)\Delta^{+}(y))-v(t,y,\mathfrak{c},s),\quad\quad v(t,y-\Delta^{-}(y),\mathfrak{c}+\mathfrak{m}Z_{-}(y)\Delta^{-}(y))-v(t,y,\mathfrak{c},s),$
represent the difference in the value function before and after a trade. If there is a sell trade, the quantity of asset $y$ in the pools jumps from $y$ to $y+\Delta^{+}(y)$, and the AMM collects fees that amount to $\mathfrak{p}\,Z_{+}(y)\Delta^{+}(y)$, i.e., the percentage of fees $\mathfrak{p}$ times the quantity traded $\Delta^{+}(y)$ times the sell exchange rate at the moment of the trade $Z_{+}(y)$. Symmetrically, when there is a buy trade, the grid for the risky asset moves from $y$ to $y-\Delta^{-}(y)$, and the AMM collects fees that amount to $\mathfrak{m}\,Z_{-}(y)\Delta^{-}(y)$. Each of the differences is multiplied by the respective intensity of order arrival.
Next, we make the ansatz $v(t,y,\mathfrak{c},s)=g(t,y,s)+\mathfrak{c}$, to obtain that $g(t,y,s)$ satisfies the HJB equation
$\frac{\partial}{\partial t}g(t,y,s)+\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial s^{2}}g(t,y,s)-P(y,s)$
$+\left(\lambda^{+}e^{k(s-Z_{+}(y))\Delta^{+}(y)}\sup_{\mathfrak{p}\in\mathbb{R}}e^{-k\mathfrak{p}Z_{+}(y)\Delta^{+}(y)}\left[g(t,y+\Delta^{+}(y),s)-g(t,y,s)+\mathfrak{p}Z_{+}(y)\Delta^{+}(y)\right]\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.2)
$+\left(\lambda^{-}e^{-k(Z_{-}(y)-s)\Delta^{-}(y)}\sup_{\mathfrak{m}\in\mathbb{R}}e^{-k\mathfrak{m}Z_{-}(y)\Delta^{-}(y)}\left[g(t,y-\Delta^{-}(y),s)-g(t,y,s)+\mathfrak{m}Z_{-}(y)\Delta^{-}(y)\right]\right)\mathbbm{1}_{\{y>\underline{y}\}}=0,$
with terminal condition $g(T,y,s)=0$. The corresponding maximizers are given by
$\mathfrak{p}^{*}(t,y)$ $:=\frac{g(t,y,s)-g(t,y+\Delta^{+}(y),s)}{Z_{+}(y)\Delta^{+}(y)}+\frac{1}{kZ_{+}(y)\Delta^{+}(y)},$ (3.3)
$\mathfrak{m}^{*}(t,y)$ $:=\frac{g(t,y,s)-g(t,y-\Delta^{-}(y),s)}{Z_{-}(y)\Delta^{-}(y)}+\frac{1}{kZ_{-}(y)\Delta^{-}(y)}.$
Plugging the maximizers in (3), we obtain
$\frac{\partial}{\partial t}g(t,y,s)+\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial s^{2}}g(t,y,s)-P(y,s)$
$\quad+\left(\frac{\lambda^{+}e^{-ks\Delta^{+}(y)-1}}{k}e^{kZ_{+}(y)\Delta^{+}(y)}e^{k(g(t,y+\Delta^{+}(y),s)-g(t,y,s))}\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.4)
$\quad+\left(\frac{\lambda^{-}e^{ks\Delta^{-}(y)-1}}{k}e^{-kZ_{-}(y)\Delta^{-}(y)}e^{k(g(t,y-\Delta^{-}(y),s)-g(t,y,s)}\right)\mathbbm{1}_{\{y>\underline{y}\}}=0.$
4 Constant oracle price
To compute the first approximate solution, we take $S_{t}=S_{0}$, which is equivalent to assuming that $\sigma=0$. In Section 5 we return to the case $\sigma>0$. With this assumption the reduced value function $g$ no longer depends on $s$ and the Hamilton-Jacobi-Bellman equation (3) simplifies to
$\frac{\partial}{\partial t}g(t,y)-P(y,s)$
$\quad+\left(\frac{\lambda^{+}e^{-ks\Delta^{+}(y)-1}}{k}e^{kZ_{+}(y)\Delta^{+}(y)}e^{k(g(t,y+\Delta^{+}(y))-g(t,y))}\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (4.1)
$\quad+\left(\frac{\lambda^{-}e^{ks\Delta^{-}(y)-1}}{k}e^{-kZ_{-}(y)\Delta^{-}(y)}e^{k(g(t,y-\Delta^{-}(y))-g(t,y)}\right)\mathbbm{1}_{\{y>\underline{y}\}}=0.$
We make the substitution $e^{kg(t,y)}=w(t,y)$ which transforms (3) into a linear system of ODEs of the form
$\frac{\partial}{\partial t}w(t,y)-kP(y,s)w(t,y)$
$+\left(\lambda^{+}e^{-ks\Delta^{+}(y)-1}e^{kZ_{+}(y)\Delta^{+}(y)}w(t,y+\Delta^{+}(y))\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (4.2)
$+\left(\lambda^{-}e^{ks\Delta^{-}(y)-1}e^{-kZ_{-}(y)\Delta^{-}(y)}w(t,y-\Delta^{-}(y))\right)\mathbbm{1}_{\{y>\underline{y}\}}=0.$
The solution to the system is summarized in the following theorem. The proof is straightforward and hence, omitted.
Theorem 4.1.
Define the matrix $\mathbf{A}:=(\mathbf{A}_{i,j})_{0\leq i\leq j\leq 2N}$ by
$\mathbf{A}_{i,j}:=\begin{cases}-kP(y^{j-N},s)&\text{ if }i=j,\\ \lambda^{+}e^{-ks\Delta^{+}(y^{j-N})-1}e^{kZ_{+}(y^{j-N})\Delta^{+}(y^{j-N})}&\text{ if }i=j-1,\\ \lambda^{-}e^{ks\Delta^{-}(y^{j-N})-1}e^{-kZ_{-}(y^{j-N})\Delta^{-}(y^{j-N})}&\text{ if }i=j+1\\ 0&\text{ otherwise,}\end{cases}$
and denote with $\mathbf{1}$ the unit vector of $\mathbb{R}^{2N}$. Define the function $w:[0,T]\times\{y^{-N},\dots,y^{N}\}\to\mathbb{R}$ by
$w(t,y^{i}):=(\exp(\mathbf{A}(T-t))\mathbf{1})_{i},$
and the function $v:[0,T]\times\{y^{-N},\dots,y^{N}\}\times\mathbb{R}_{+}\to\mathbb{R}$ as
$v(t,y^{i},\mathfrak{c})=\mathfrak{c}+\frac{1}{k}\log(w(t,y^{i})).$
Then $v$ solves the Hamilton-Jacobi-Bellman equation
$\frac{\partial}{\partial t}v(t,y^{i},\mathfrak{c})-P(y^{i},s)+$
$\left(\lambda^{+}e^{-ks\Delta^{+}(y^{i})}\sup_{\mathfrak{p}\in\mathbb{R}}e^{kZ_{+}(\mathfrak{p},y^{i})\Delta^{+}(y^{i})}\left[v(t,y^{i+1},\mathfrak{c}+\mathfrak{p}Z_{+}(y^{i})\Delta^{+}(y^{i}))-v(t,y^{i},\mathfrak{c})\right]\right)\mathbbm{1}_{\{y^{i}<y^{N}\}}$
$\left(\lambda^{-}e^{ks\Delta^{-}(y^{i})}\sup_{\mathfrak{m}\in\mathbb{R}}e^{-kZ_{-}(\mathfrak{m},y^{i})\Delta^{-}(y^{i})}\left[v(t,y^{i-1},\mathfrak{c}+\mathfrak{m}Z_{-}(y^{i})\Delta^{-}(y^{i}))-v(t,y^{i},\mathfrak{c})\right]\right)\mathbbm{1}_{\{y^{i}>y^{-N}\}}=0,$
with boundary condition $v(T,y,\mathfrak{c})=\mathfrak{c}$. Moreover, the corresponding optimizers are independent of $\mathfrak{c}$ and satisfy
$\mathfrak{p}^{*}(t,y^{i})$ $:=\frac{1}{kZ_{+}(y^{i})\Delta^{+}(y^{i})}\left(1+\log\left(\frac{w(t,y^{i})}{w(t,y^{i+1})}\right)\right),$ (4.3)
$\mathfrak{m}^{*}(t,y^{i})$ $:=\frac{1}{kZ_{-}(y^{i})\Delta^{-}(y^{i})}\left(1+\log\left(\frac{w(t,y^{i})}{w(t,y^{i-1})}\right)\right),$ (4.4)
for $y^{i}\in\{y^{-N}=\underline{y},\dots,y^{0}=y_{0},\dots,y^{N}=\overline{y}\}$.
4.1 Optimal fee structure
We perform numerical simulations of the results found in Theorem 4.1. We assume that the AMM is a CPM, where the trading function $f:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is $f(x,y):=xy$ and the level function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is $\varphi(y):=0pt/y$. We assume that the the grid for $y$ is such that after every trade the exchange rate $Z(y)$ moves by $0.1$. More precisely, the grid for the risky asset is in turn given by the formula
$y^{i}:=\sqrt{\frac{0pt}{\left(\frac{0pt}{y_{0}}-0.1\,i\right)}},\quad\quad\text{ for }i\in\{-20,\dots,20\}.$
We take the time horizon $T=1$, the two baseline intensities $\lambda^{-}=\lambda^{+}=50$, the rate of exponential decay $k=2$, the oracle price $S_{0}=100$, and the penalty function $P(y,s)=\phi(Z(y)-s)^{2}$. Moreover, we assume that the initial conditions of the pool are $0pt=10^{8}$ and $y_{0}=1000$. Figure 2(a) shows the optimal fees in the case where the penalization constant $\phi=0$.
From the plot in Figure 2(a) we see that the AMM charges fees in two different regimes. When the quantity of asset $y$ is below the value of $y_{0}$ it is more profitable to sell inside the AMM with respect to the oracle price, hence the AMM will charge high fees to penalize arbitrageurs. At the same time it will charge low fees to buy, even negative ones, in order to encourage noise traders to trade inside the venue and increase the volatility in the price. Symmetrically, the analogous happens when the quantity is above the value of $y_{0}$. Figure 2(b) shows that if the quantity of asset $Y$ stays in a small enough neighbourhood of $y_{0}$, a linear approximation of the fees is a good enough approximation. Below we show that the linear approximation generates revenues close to the optimal strategy.
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
(b) Linear fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
The behaviour of the fees is highly dependent on the value of the parameter $k$. Indeed, as the value of the parameter $k$ gets smaller, the probability of arrival becomes less sensitive to price changes. As a response to this, the AMM prioritizes pushing the price as close to the oracle price as possible to get away from the boundaries where it does not generate revenues. We can observe this phenomenon in Figure 3(a) and Figure 3(b) for $k=1$ and $k=0.25$, respectively.
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) multiplied by $k$ at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool when $k=1$.
(b) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) multiplied by $k$ at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool when $k=0.25$.
The following corollary gives a rigorous mathematical characterization of this behaviour. The proof is straightforward and hence, omitted.
Corollary 4.2.
Define the matrix $\mathbf{A}:=(\mathbf{A}_{i,j})_{0\leq i\leq j\leq 2N}$ by
$\mathbf{A}_{i,j}:=\begin{cases}\lambda^{+}e^{-1}&\text{ if }i=j-1,\\ \lambda^{-}e^{-1}&\text{ if }i=j+1,\\ 0&\text{ otherwise,}\end{cases}$
and denote with $\mathbf{1}$ the unit vector of $\mathbb{R}^{2N}$. Define the function $w:[0,T]\times\{y^{-N},\dots,y^{N}\}\to\mathbb{R}$ by
$w(t,y^{i}):=(\exp(\mathbf{A}(T-t))\mathbf{1})_{i},.$
Let $\mathfrak{p}^{*}$ and $\mathfrak{m}^{*}$ be the optimal fees from Theorem 4.1 and define the quantities
$\mathfrak{p}^{*}_{0}(t,y^{i})$ $:=\frac{1}{Z_{+}(y^{i})\Delta^{+}(y^{i})}\left(1+\log\left(\frac{w(t,y^{i})}{w(t,y^{i+1})}\right)\right)$ (4.5)
$\mathfrak{m}^{*}_{0}(t,y^{i})$ $:=\frac{1}{Z_{-}(y^{i})\Delta^{-}(y^{i})}\left(1+\log\left(\frac{w(t,y^{i})}{w(t,y^{i-1})}\right)\right).$ (4.6)
We have that
$\lim_{k\to 0}k\mathfrak{p}^{*}(t,y^{i})=\mathfrak{p}^{*}_{0}(t,y^{i})\quad\quad\text{ and }\quad\quad\lim_{k\to 0}k\mathfrak{m}^{*}(t,y^{i})=\mathfrak{m}^{*}_{0}(t,y^{i}),$ (4.7)
for every $t\in[0,T]$ and $y^{i}\in\{y^{-N},\dots,y^{N}\}$.
The next plots illustrates the corollary. More precisely, we compare $k*\mathfrak{p}^{*}$ and $k*\mathfrak{m}^{*}$ for $k=0.1$ (left figure) with the limit optimal fees $\mathfrak{p}^{*}_{0}$ and $\mathfrak{m}^{*}_{0}$ (right figure) in (4.2).
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) multiplied by $k$ at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool when $k=0.5$.
(b) Optimal fees for selling $\mathfrak{p}_{0}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}_{0}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
Next we discuss how the optimal fees behave as functions of $\phi$. If we increase $\phi$, the AMM tries to keep prices aligned to the external price by pushing quantities of asset $Y$ away from the boundaries. To this end, it charges low fees to sell and high fees to buy when the price is less than the oracle price and vice versa when the price is high. Figure 5(a) and Figure 5(b) show the fees as a function of quantity and penalization constant when $t=0.5$, for $k=2$ and $k=0.1$ respectively. The curves corresponding to $\phi=0$ are the ones in Figure 2(a) and Figure 4(a) respectively .
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool and of $\phi$ when $k=2$.
(b) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ and of $\phi$ in the pool when $k=0.1$.
As the penalty parameter $\phi$ increases, it alters both the shape and slope of the optimal fees. When $\phi$ becomes sufficiently large, the penalty term dominates the allocation of the fees, with alignment of quotes becoming a secondary effect. The following plots show the behaviour of the fees as a function of time ($x$ axis) and quantity in the pool (colorbar) for $k=2$ and $k=0.1$. The effect present in the Figure 4(a) can be seen also in Figure 6(b).
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ (colorbar) in the pool and of time when $k=2$.
(b) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ (colorbar) and of time in the pool when $k=0.1$.
Lastly we compare our optimal strategy with other different strategies:
  • $(i)$
    the linear strategy described in Figure 2(b), where at each time with linearised $\mathfrak{p}^{*}(t,y)$, $\mathfrak{m}^{*}(t,y)$ in a neighbourhood of $y_{0}$, and
  • $(ii)$
    the constant strategy where the fees are constant for every time $t$ and every quantity $y$. The constant $c$ is chosen as the average of the optimal two fees at $t=0.5$ for $y=y_{0}$, i.e.,
    $c=\frac{\mathfrak{p}^{*}(0.5,y_{0})+\mathfrak{m}^{*}(0.5,y_{0})}{2}.$
We consider the following parameters $0pt=10^{8}$, $y_{0}=1000$, $\phi=0$, $\underline{y}=980$, $\overline{y}=1020$, $T=1$ and we run simulations for different values of $\lambda^{-}$, $\lambda^{+}$ and $k$. We carry out 100,000 simulations and we discretise $[0,T]$ in $1,000$ timesteps. [^5] Our code is publicly available at https://github.com/leonardobaggiani/amm-fees.git.
We observe that the revenues are decreasing in the values of $k$ and increasing in the value of $\lambda^{+}$ and $\lambda^{-}$. The latter is because as $\lambda^{+}$ and $\lambda^{-}$ increase the number of orders increase and hence the revenue increases, similarly, as $k$ decreases, the arbitrageurs react less to mispricing and simultaneously, the optimal fees increase and thus the AMM collects more fees.
In the following table, the column fees shows the revenue from collecting fees, the column sell shows the number of sell orders, the column buy shows the number of buy orders, and the column QV shows the quadratic variation of the instantaneous exchange rate $Z$ defined in Equation 1.
$\lambda^{+}=\lambda^{-}=100$ $\lambda^{+}=\lambda^{-}=150$
fees sell buy QV fees sell buy QV
$k=2$
Optimal 35.55 35.89 35.91 0.69 52.97 53.45 53.47 1.01
Linear 35.55 35.88 35.91 0.69 52.97 53.44 53.48 1.01
Constant 35.16 35.15 35.16 0.68 52.26 52.26 52.26 0.99
$k=1$
Optimal 71.46 35.92 35.94 0.69 106.38 53.47 53.50 1.01
Linear 71.46 35.91 35.95 0.69 106.38 53.46 53.51 1.01
Constant 71.22 35.60 35.62 0.69 105.90 52.94 52.96 1.00
We find that optimal strategy outperforms the constant strategy and the linear approximation of the fees has a performance that is indistinguishable from that of the optimal strategy for the above parameters. Thus, a linear fee structure is suitable when designing the fees to charge in these venues.
5 Quadratic approximation with stochastic oracle price
For the second approximate solution we return to the original HJB equation in (3). Here, we make a quadratic approximation of the exponential terms:
$e^{k(-s\Delta^{+}(y)+Z_{+}(y)\Delta^{+}(y)+g(t,y+\Delta^{+}(y),s)-g(t,y,s))}\approx 1$ $+k(-s\Delta^{+}(y)+Z_{+}(y)\Delta^{+}(y)+g(t,y+\Delta^{+}(y),s)-g(t,y,s))$
$+\frac{1}{2}[k(-s\Delta^{+}(y)+Z_{+}(y)\Delta^{+}(y)+g(t,y+\Delta^{+}(y),s)-g(t,y,s))]^{2},$
$e^{k(s\Delta^{-}(y)-Z_{-}(y)\Delta^{-}(y)+g(t,y-\Delta^{-}(y),s)-g(t,y,s))}\approx 1$ $+k(s\Delta^{-}(y)-Z_{-}(y)\Delta^{-}(y)+g(t,y-\Delta^{-}(y),s)-g(t,y,s))$
$+\frac{1}{2}[k(s\Delta^{-}(y)-Z_{-}(y)\Delta^{-}(y)+g(t,y-\Delta^{-}(y),s)-g(t,y,s))]^{2}.$
The HJB in (3) becomes
$\frac{\partial}{\partial t}g(t,y,s)$ $+\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial s^{2}}g(t,y,s)-\phi(Z(y)-s)^{2}+$ (5.1)
$\frac{e^{-1}\lambda^{+}}{k}[1-k(s\Delta^{+}(y)+Z_{+}(y)\Delta^{+}(y)+g(t,y+\Delta^{+}(y),s)-g(t,y,s))$
$+\frac{k^{2}}{2}(-s\Delta^{+}(y)+Z_{+}(y)\Delta^{+}(y)+g(t,y+\Delta^{+}(y),s)-g(t,y,s))^{2}]$
$\frac{e^{-1}\lambda^{-}}{k}[1+k(s\Delta^{-}(y)-Z_{-}(y)\Delta^{-}(y)+g(t,y-\Delta^{-}(y),s)-g(t,y,s))$
$+\frac{k^{2}}{2}(s\Delta^{-}(y)-Z_{-}(y)\Delta^{-}(y)+g(t,y-\Delta^{-}(y),s)-g(t,y,s))^{2}]=0.$
We assume that $\Delta^{-}(y)$ and $\Delta^{+}(y)$ are constant equal to $\delta^{-}$ and $\delta^{+}$, respectively.
Remark 5.1.
The assumption that $\Delta^{-}(y)$ and $\Delta^{+}(y)$ are constant is necessary if one wishes to carry out a quadratic expansion of the exponential terms. If we employ a linear approximation of the exponential terms, we could work with the case where $\Delta^{-}(y)$ and $\Delta^{+}(y)$ are linear in $y$.
Lastly, we use a linear approximation to the exchange rates
$Z(y)$ $=\frac{0pt}{y^{2}}\approx\frac{0pt}{y_{0}^{2}}-2(y-y_{0})\frac{0pt}{y_{0}^{3}},$
$Z_{+}(y)\delta^{+}$ $=\frac{0pt}{y^{2}+\delta^{+}y}\approx\frac{0pt}{y_{0}^{2}+\delta^{+}y_{0}}-(y-y_{0})\frac{0pt(2y_{0}+\delta^{+})}{(y_{0}^{2}+\delta^{+}y_{0})^{2}},$
$Z_{-}(y)\delta^{-}$ $=\frac{0pt}{y^{2}-\delta^{-}y}\approx\frac{0pt}{y_{0}^{2}-\delta^{-}y_{0}}-(y-y_{0})\frac{0pt(2y_{0}-\delta^{-})}{(y_{0}^{2}-\delta^{-}y_{0})^{2}}.$
In order to find a solution for the equation (5.1) we assume that $g(t,y,s)$ is a quadratic polynomial in $y$ of the form
$g(t,y,s)=y^{2}A(t)+yB(t,s)+C(t,s),$
with $A(T)=B(T,s)=C(T,s)=0$. Proposition (5.2) formalises the solution to the HJB equation.
Theorem 5.2.
There exists constants [^6] The full definition of the constants $\{\psi_{i}\}_{1\leq i\leq 26}$ is publicly available in the GitHub of the project. $\{\psi_{i}\}_{1\leq i\leq 26}\subset\mathbb{R}$ that define the unique solution to the HJB equation (5.1) as follows. Let $A:[0,T]\to\mathbb{R}$ be the solution to the Riccati ODE
$\psi_{0}+\psi_{1}A(t)+\psi_{2}A^{2}(t)+A^{\prime}(t),\quad\quad A(T)=0.$
Let $b_{0}$, $b_{1}:[0,T]\to\mathbb{R}$ be the unique solutions to the system
$\begin{cases}\psi_{3}+\psi_{4}A(t)+\psi_{5}A^{2}(t)+\psi_{6}A(t)b_{0}(t)+\psi_{7}b_{0}(t)+b_{0}^{{}^{\prime}}(t)=0\\ \psi_{8}+\psi_{9}A(t)+\psi_{10}A(t)b_{1}(t)+\psi_{11}b_{1}(t)+b_{1}^{{}^{\prime}}(t)=0,\end{cases}$
with terminal conditions $b_{0}(T)=b_{1}(T)=0$. Let $c_{0}$, $c_{1}$, $c_{2}:[0,T]\to\mathbb{R}$ be the unique solutions to the system
$\begin{cases}\psi_{12}+\psi_{13}A(t)+\psi_{14}A^{2}(t)+\psi_{15}A(t)b_{0}(t)+\psi_{16}b_{0}(t)+\psi_{17}b_{0}^{2}(t)+\sigma^{2}c_{2}(t)+c_{0}^{{}^{\prime}}(t)=0,\\ \psi_{18}+\psi_{19}A(t)+\psi_{20}A(t)b_{1}(t)+\psi_{21}b_{0}(t)+\psi_{22}b_{0}(t)b_{1}(t)+\psi_{23}b_{1}(t)+c_{1}^{{}^{\prime}}(t)=0,\\ \psi_{24}+\psi_{25}b_{1}(t)+\psi_{26}b_{1}^{2}(t)+c_{2}^{{}^{\prime}}(t)=0,\end{cases}$
with terminal conditions $c_{0}(T)=c_{1}(T)=c_{2}(T)=0$. Define $\hat{B},\hat{C}:[0,T]\times\mathbb{R}\to\mathbb{R}$ as
$\hat{B}(t,s)$ $=sb_{1}(t)+b_{0}(t),$
$\hat{C}(t,s)$ $=s^{2}c_{2}(t)+sc_{1}(t)+c_{0}(t).$
Then, the function
$g(t,y,s)=y^{2}\hat{A}(t)+y\hat{B}(t,s)+\hat{C}(t,s),$
is a solution to the HJB equation (5.1).
With the above results, the optimal fees from (3.3) become
$\mathfrak{p}^{*}(t,y)$ $:=-\frac{(2y+\delta^{+})\hat{A}(t)+\hat{B}(t,s)}{Z_{+}(y)}+\frac{1}{kZ_{+}(y)\delta^{+}},$ (5.2)
$\mathfrak{m}^{*}(t,y)$ $:=-\frac{(-2y+\delta^{-})\hat{A}(t)-\hat{B}(t,s)}{Z_{-}(y)}+\frac{1}{kZ_{-}(y)\delta^{-}}.$
Observe that the fees depend linearly on $y$ and $s$ and they do not depend on $\sigma$.
5.1 Optimal fee structure: dynamic oracle price with second order approximation
We perform numerical simulations of the results found in Theorem 5.2. In order to make it consistent with the hypothesis of the theorem, we assume that the traded amounts are $\Delta^{+}(y)=\delta^{+}=0.5$ and $\Delta^{-}(y)=\delta^{-}=0.5$. We fix an initial value of $y_{0}=1000$ which gives the following grid for $y$ as
$y^{i}=(1000+i),\quad\quad\text{ for }i\in\{-20,-19.5,\dots,19.5,20\}.$
We take the time horizon $T=1$, the two baseline intensities $\lambda^{-}=\lambda^{+}=50$, the rate of exponential decay $k=2$, the oracle price $S_{t}=S_{0}+\sigma W_{t}$ with $S_{0}=100$, $\sigma=0.2$ and $\phi=0$. As before, we assume that the depth of the pool is $0pt=10^{8}$.
The next figures show the plot for the optimal fees of (5.2) and the comparison with the optimal fees obtained with the first approximation. It is important to point out that the fees in (5.2) are no longer skewed at the boundaries.
(a) Optimal fees (second approximation Section 5) for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
(b) Optimal fees (first approximation Section 4) for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
The plot in Figure 7(a) confirm the insights we obtained from Figure 2(a), namely that there are two distinct regimes for the fee structure: one in which arbitrageurs are penalised and one that aim to increase volatility and noise trading. Next, we study the case when $k\to 0$. Here, the second order approximation does not work well for smaller values of the parameters $k$. Indeed, since the fees are independent of $\sigma$ we would expect that as $k\to 0$ the fees in (5.2) should resemble the one in Figure 4(b), however, the next two figures show that this is not the case. The explanation for this phenomenon comes from the fact that when $k\to 0$ then $A(t)\to 0$ and $B(t)\to\psi_{3}t$, hence the following holds.
Corollary 5.3.
Let $\mathfrak{p}^{*}(t,y)$ and $\mathfrak{m}^{*}(t,y)$ be the optimal fees from Theorem 5.2 and define the quantities
$\mathfrak{p}^{*}_{0}(t,y^{i})$ $:=\frac{1}{Z_{+}(y^{i})\delta^{+}}$ (5.3)
$\mathfrak{m}^{*}_{0}(t,y^{i})$ $:=\frac{1}{Z_{-}(y^{i})\delta^{-}}.$ (5.4)
Then we have that
$\lim_{k\to 0}k\mathfrak{p}^{*}(t,y^{i})=\mathfrak{p}^{*}_{0}(t,y^{i})\quad\quad\text{ and }\quad\quad\lim_{k\to 0}k\mathfrak{m}^{*}(t,y^{i})=\mathfrak{m}^{*}_{0}(t,y^{i}),$ (5.5)
for every $t\in[0,T]$ and $y^{i}\in\{y^{-N},\dots,y^{N}\}$.
Figures 8(a) and 8(b) illustrate the limiting behaviour described in Corollary (5.3).
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) for $k=0.25$ at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
(b) Optimal fees for selling $\mathfrak{p}_{0}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}_{0}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool.
The next two figures show how the optimal fees behave as functions of $\phi$ (left) and as a function of $S_{t}$ (right). Similarly to the first approximation there are two different regimes, one for $\phi=0$ and one for $\phi\neq 0$. Indeed, when $\phi\neq 0$ the AMM tries to push the price away from the boundaries. Hence, when the quantity of asset $y$ is close to $\underline{y}$ it will charge high fees to buy and low (or even negative) fees to sell and when the quantity of asset $y$ is close to $\overline{y}$ we get the symmetric effect. As expected from the expression in (5.2), the dependence on $S_{t}$ is linear, furthermore, $\mathfrak{p}$ decreases when $S_{t}$ increases and $\mathfrak{m}$ decreases as $S_{t}$ increases.
(a) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool and $\phi$ (colorbar).
(b) Optimal fees for selling $\mathfrak{p}^{*}(t,y_{t-})$ (solid line) and for buying $\mathfrak{m}^{*}(t,y_{t-})$ (dashed line) at time $t=0.5$ as a function of the quantity of asset $Y$ in the pool and $s$ (colorbar).
We now compare the value functions obtained from the two approximations. If $\sigma=0$ then the value function $v(t,y,\mathfrak{c})$ (right) from Theorem 4.1 is the true value function and we expect it to be close to the function $g(t,y,s)$ (left) in the case of $\sigma=0$. The following figures shows the comparison between the plots for $g$ and $v$ depending from time and the quantity of asset $Y$. The interval $[0,1]$ has been discretized in $1000$ time steps.
(a) Value function first approximation.
(b) Second approximation when $\sigma=0$.
(c) Difference between the two approximations.
Finally we compare the optimal strategy from the second-order approximation in (5.2) (Optimal SA) to the following strategies:
  • $(i)$
    the optimal strategy from the first approximation (Optimal FA) implemented with a stochastic $S_{t}$ instead of $S_{0}$,
  • $(ii)$
    the constant strategy where the fees are constant for every time $t$ and every quantity $y$. The constant $c$ is chosen as the average of the optimal two fees at $t=0.5$ for $y=y_{0}$, i.e.,
    $c=\frac{\mathfrak{p}^{*}(0.5,y_{0})+\mathfrak{m}^{*}(0.5,y_{0})}{2}.$
We use the following parameters: $0pt=10^{8}$, $y_{0}=1000$, $\phi=0$, $\underline{y}=990$, $\overline{y}=1010$, $T=1$ and we run simulations for different $\lambda^{-}$, $\lambda^{+}$ and $k$. We carry out 100,000 simulations and we discretise $[0,T]$ in $1,000$ timesteps.
In the following table the column fees shows the revenue from collecting fees, the column sell shows the number of sell orders, the column buy shows the number of buy orders and the column QV shows the quadratic variation of the instantaneous exchange rate $Z$ defined in Equation 1.
$\lambda^{+}=\lambda^{-}=100$ $\lambda^{+}=\lambda^{-}=150$
fees sell buy QV fees sell buy QV
$k=2$
Optimal SA 35.62 35.90 35.91 0.69 53.05 51.60 51.60 0.98
Optimal FA 35.61 34.79 34.76 0.67 53.02 53.45 53.45 1.01
Constant 35.18 35.18 35.19 0.68 52.29 52.28 52.29 0.99
$k=1$
Optimal SA 71.49 35.92 35.93 0.69 106.42 53.46 53.47 1.00
Optimal FA 71.51 35.47 35.44 0.68 106.38 53.46 53.51 1.01
Constant 71.23 35.60 35.61 0.69 105.90 52.91 52.91 1.00
As before, we find that either approximation yields a performance that is almost indistinguishable from the other, and both outperform constant fees substantially. This supports the usage of linear dynamic fees when designing fee structures in AMMs.
6 Conclusion
In this paper, we study how an AMM maximizes its revenue by setting optimal dynamic trading fees. We find that the optimal fee schedule serves two distinct purposes. On the one hand, by monitoring external prices, the AMM raises fees strategically to penalize arbitrageurs. On the other hand, it lowers fees to increase price volatility, which attracts noise traders and stimulates trading volume. This dual approach allows the AMM to balance the trade-off between deterring arbitrageurs and encouraging liquidity taking through noise order flow.
Although in this paper we focus on a time-horizon that is small enough so that liquidity remains constant, future work will look at the case of non-static liquidity; see e.g., Aqsha et al. (2025); Bergault et al. (2024a). Similarly, one could also look at online learning algorithms (e.g., adaptation and learning with neural contextual bandits or reinforcement learning, as in Duran-Martin et al. (2022); Spooner et al. (2018)) to automate the fees that AMMs charge. Competition between pools is another promising area of future work; see e.g., Boyce et al. (2025); Guo et al. (2024) for studies in limit order books.

### v2

@@
 We find that there are two distinct fee regimes: one in which the AMM imposes higher fees to deter arbitrageurs, and another where fees are lowered to increase volatility and attract noise traders.
-Our results also show that dynamic fees that are linear in inventory and are sensitive to changes in the external price are a good approximation of the optimal fee structure and thus constitute suitable candidates when designing fees for AMMs.
+Our results also show that dynamic fees that are linear in inventory and that track changes in the external price are a good approximation of the optimal fee structure and thus constitute suitable candidates when designing fees for AMMs.
 Keywords: decentralized finance, automated market makers, optimal fees, arbitrageurs, noise trading.
@@
 (2021), Angeris and Chitra (2020), Angeris et al.
-(2021), and Angeris et al.
+(2021), Cartea et al.
+(2022) and Angeris et al.
 (2023) first explain how these exchanges operate and introduce the problem of liquidity provision in AMMs.
 As shown by Cartea et al.
-(2023, 2024b) and Milionis et al.
+(2023, 2024b), Milionis et al.
+(2022), and Fukasawa et al.
 (2025), liquidity provision in the absence of fees leads to losses for liquidity providers.
-This issue, known as impermanent loss (IL), occurs because liquidity provision in an AMM is passive; when the prices of risky assets move, the liquidity providers trade at worse than market prices.
-To address the impermanent loss problem, a significant part of the research has focused on improving the design of AMMs to make liquidity provision less exposed to arbitrageurs.
-Some works in this direction are Adams et al.
-(2025), where they set up an auction-based AMM in which LPs are rewarded by a so-called manager of the pool who can profit from arbitrage opportunities; Goyal et al.
-(2023), where they translate LPs’ beliefs about future asset valuations into an optimal choice of a trading function; Cartea et al.
-(2024a), where they introduce a more general version of constant function market makers (CFMs) called decentralized liquidity pools (DLPs), in which liquidity providers prevent predictable losses by choosing particular quote functions, impact functions, and monitoring external information; and Bergault et al.
-(2024b), where they design an oracle-based AMM in which price discovery does not rely solely on liquidity takers, and the pricing function incorporates external information about the current market price of the risky asset.
-Recently Capponi and Zhu (2025) addressed the problem of IL by determining an optimal exit strategy for the liquidity provider.
-Another interesting contribution is the work of Fukasawa et al.
-(2023), where the authors hedge impermanent loss in a constant function market maker using a weighted variance swap.
+There are two well-understood types of losses: loss-versus-rebalancing (LVR) and impermanent loss (IL).
+Milionis et al.
+(2024) describe LVR as
+the loss incurred by LPs relative to a trading strategy (the “rebalancing strategy”) that holds the same risky positions as the pool, but that trades at market prices rather than AMM prices.
+On the other hand, IL is the difference in value between holding the assets inside the pool and holding the assets outside of the pool.
+The name impermanent is due to the fact that if the market condition reverse, the losses disappear.
+To address the IL issue, the literature has generally taken two main approaches.
+The first is to redesign automated market makers (AMMs) to better account for spot price volatility.
+For instance, Bergault et al.
+(2024b) propose an oracle-based AMM, where price discovery is not determined solely by liquidity takers.
+Instead, their pricing function integrates external information about the current market price of the risky asset.
+The second approach focuses on finding trading strategies that allow liquidity providers to adapt to changing market conditions.
+Capponi and Zhu (2025) tackle the issue of IL by determining an optimal exit strategy for liquidity providers.
+Fukasawa et al.
+(2023) address IL in constant function market makers by employing a weighted variance swap for hedging, and in a follow up study, Fukasawa et al.
+(2024) demonstrate a super-hedging strategy for IL that holds regardless of the oracle price dynamics.
+Most recently, Sepp et al.
+(2024) show that IL can be fully hedged using European put and call options.
+To find solutions for the loss-versus-rebalancing problem, much of the research has centered on improving AMM designs to make liquidity provision less vulnerable to arbitrageurs.
+Several notable works in this direction include Adams et al.
+(2025), who introduce an auction-based AMM in which a pool “manager” captures arbitrage profits and rewards liquidity providers accordingly; Goyal et al.
+(2023), who convert LPs’ beliefs about future asset values into optimal trading functions; and Cartea et al.
+(2024a), who generalize constant function market makers into decentralized liquidity pools (DLPs), enabling LPs to mitigate loss-versus-rebalancing by selecting suitable quote and impact functions and by monitoring external information.
+Another notable contribution is the work by Nadkarni et al.
+(2024), in which the authors investigate the problem of finding an optimal dynamic trading curve with the focus is specifically on developing a strategy that effectively hedges against loss-versus-rebalancing.
 As shown by Milionis et al.
-(2024), impermanent loss can be partially or fully hedged through the collection of fees.
-The problem of determining optimal fees for AMMs has been an important issue since the early stages of research in this area.
+(2024), loss-versus-rebalancing can be partially or fully hedged through the collection of fees.
+Thus, the problem of determining optimal fees for AMMs has been an important issue since the early stages of research in this area.
 Evans et al.
 (2021) focused on how to find optimal fees for geometric mean market makers (G3Ms) while Fritsch (2021) studied how to determine optimal fees in a setting of multiple liquidity pools competing with each other for trade volume.
-However, both works only address the problem of finding constant static fees.
-Sabate-Vidales and Šiška (2023) address the impermanent loss problem by setting dynamic fees although no optimality is established.
+He et al.
+(2024) determined the optimal fee structure for a liquidity provider who allocates their capital between an automated market maker (AMM) and centralized trading venues.
+However, these works only address the problem of finding constant static fees.
+Sabate-Vidales and Šiška (2023) address the loss-versus-rebalancing problem by setting dynamic fees although no optimality is established.
 Hasbrouck et al.
@@
 $\frac{\partial}{\partial t}g(t,y,s)+\frac{\sigma^{2}}{2}\frac{\partial^{2}}{\partial s^{2}}g(t,y,s)-P(y,s)$
-$+\left(\lambda^{+}e^{k(s-Z_{+}(y))\Delta^{+}(y)}\sup_{\mathfrak{p}\in\mathbb{R}}e^{-k\mathfrak{p}Z_{+}(y)\Delta^{+}(y)}\left[g(t,y+\Delta^{+}(y),s)-g(t,y,s)+\mathfrak{p}Z_{+}(y)\Delta^{+}(y)\right]\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.2)
+$+\left(\lambda^{+}e^{k(Z_{+}(y)-s)\Delta^{+}(y)}\sup_{\mathfrak{p}\in\mathbb{R}}e^{-k\mathfrak{p}Z_{+}(y)\Delta^{+}(y)}\left[g(t,y+\Delta^{+}(y),s)-g(t,y,s)+\mathfrak{p}Z_{+}(y)\Delta^{+}(y)\right]\right)\mathbbm{1}_{\{y<\overline{y}\}}$ (3.2)
 $+\left(\lambda^{-}e^{-k(Z_{-}(y)-s)\Delta^{-}(y)}\sup_{\mathfrak{m}\in\mathbb{R}}e^{-k\mathfrak{m}Z_{-}(y)\Delta^{-}(y)}\left[g(t,y-\Delta^{-}(y),s)-g(t,y,s)+\mathfrak{m}Z_{-}(y)\Delta^{-}(y)\right]\right)\mathbbm{1}_{\{y>\underline{y}\}}=0,$
@@
 (2022); Spooner et al.
-(2018)) to automate the fees that AMMs charge.
+(2018)) to automate the fees that AMMs charge, similar to what has been done in Nadkarni et al.
+(2023).
 Competition between pools is another promising area of future work; see e.g., Boyce et al.